OpenStudy (anonymous):

Only have about 3 questions today. And I'm pretty sure I know the answers, I just want to make sure they're correct :) @jigglypuff314

OpenStudy (anonymous):

Im pretty sure p = 3?

OpenStudy (anonymous):

6(2p-3) = 4(4.5)

jigglypuff314 (jigglypuff314):

mmm @jdoe0001 is better at this stuff :P idk

jigglypuff314 (jigglypuff314):

if it is that, then 3 is correct :)

OpenStudy (anonymous):

Can @jdoe0001 Help? I want to be sure im doing it correctly :/

OpenStudy (anonymous):

I don't get this one :/

jigglypuff314 (jigglypuff314):

http://www.mathopenref.com/secantsintersecting.html p(p+q) = w(w+v)

OpenStudy (anonymous):

Oh yeah i know, I got it :)

OpenStudy (anonymous):

Thanks for the help. That's all xD Good day! ;o